OpenStudy (help108):

Medal FAN!!! What is the solution to the equation 4x + 2(x − 3) = 3x + x − 12?

OpenStudy (help108):

@AloneS

OpenStudy (help108):

help anyone

OpenStudy (help108):

@AloneS

OpenStudy (help108):

Alone you were typing now not anymore.

alones (alones):

4x+2(x−3)=3x+x−12 simplify both sides of the equation. 4x+2(x−3)=3x+x−12 6x−6=4x−12 then subtract 4x from both sides. 6x−6−4x=4x−12−4x 2x−6=−12 add 6 to both sides. 2x−6+6=−12+6 2x=−6 the last stpe is to divide both sides by 2. 2x/2=−6/2 x=−3

OpenStudy (help108):

thanks